OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 2) = 4x. negative 2 over 5 negative 5 over 2 2 over 5 5

OpenStudy (anonymous):

is that \[\frac{ 2 }{ 3(x-2) }=4x ?\]

OpenStudy (anonymous):

2/3 (x-2) = 4x

OpenStudy (anonymous):

ok, start by getting rid of parenthesis

OpenStudy (anonymous):

i honestly am not good at math -.- how would i do that

OpenStudy (anonymous):

unless, you just want the answer....if that's the case \[-\frac{2}{5}\]

OpenStudy (anonymous):

thanks :)